OpenStudy (anonymous):

Solve for x in the equation ln(2x – 1) – ln(x -1) = ln 5.

OpenStudy (anonymous):

you would have to know that \[\ln A -\ln B = \ln \frac { A }{ B }\]

OpenStudy (anonymous):

so using that you get \[\ln \frac{ 2x-1 }{ x-1 } = \ln 5\]

OpenStudy (anonymous):

NICE PRACTICE @kausarsalley

OpenStudy (anonymous):

so now using the definition of log, which says \[\log _ba=x\] is the same as writing \[b ^{x}=a\]

OpenStudy (anonymous):

sorry for the late reply......having problems

OpenStudy (anonymous):

\[\frac{ 2x-1 }{ x-1 }=e ^{\ln 5}\]

OpenStudy (anonymous):

can you continue from there??? or should i further explain??

OpenStudy (anonymous):

Can you explain more?

OpenStudy (anonymous):

\[2x-1=e ^{\ln 5}(x-1)\] \[2x-1=e ^{\ln 5}x-e ^{\ln5}\]

OpenStudy (anonymous):

\[e ^{\ln 5}x-2x=e ^{\ln 5}-1\] \[5x-2x=5-1\] \[3x=4\] \[x=\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

@sfb